You are considering purchasing a CNC machine which costs ​$130,000. This machine will have an estimated service life of 99 years with a net​ after-tax salvage value of ​$13,000. Its annual​after-tax operating and maintenance costs are estimated to be ​$57,000. To expect an 20​% rate of return on​ investment, what would be the required minimum annual​ after-tax revenues?

The required minimum annual​ after-tax revenues would be ​$____thousand. ​(Round to one decimal​ place.)
Answer #1

Let the revenue per year be X

Present value of future cash flow=present value value of annual cash flow+present value of salvage value

PW (i) = [D+A(P/A,i,N)+F(P/F,i,N)

PW(0.20) = -130000+(X-57000)(P/A,0.20,99)+13000(P/F,0.20,99)]

PW(0.20) = -130000+(X-57000)(5)+13000(0.0000000145)

PW = -130000+5X-285000+0.0001885

5X=414999.99

X=$82999.9 is the required minimum annual after tax revenues